# ReALM: Reference Resolution As Language Modeling

### v1

@@ 6.1 Analysis
GPT-4 $\approx$ ReaLM $\gg$ MARRS for new use-cases: As a case study, we explore zero-shot performance of this model on an unseen domain: Alarms (we show a sample data point in Appendix Table 11). The last row in Table 3 compares the performance of all approaches and baselines on this unseen test set. We find that all of the LLM-based approaches outperform the FT model for this test set. Among the two, we find that the performance of ReaLM and GPT-4 are very similar for the unseen domain. Additionally, table 4 shows completely new experiences enabled by ReaLM due to the LLM’s superior ability to perform complex understanding of natural language.
ReaLM $>$ GPT-4 for domain-specific queries We find that due to finetuning on user requests, ReaLM is able to understand more domain-specific questions. For instance, consider Table 4. GPT-4 incorrectly assumes the reference to be about only a setting, whereas the ground truth consists of a home automation device in the background as well, and GPT-4 lacks the domain knowledge to be able to recognise that the device would also be relevant to this reference. ReaLM, in contrast, doesn’t suffer from this due to being trained on domain-specific data.
@@ 7 Conclusion and Future Work
In this work, we demonstrate how large language models can be used to perform reference resolution. We do this by encoding entity candidates as natural text; critically, we demonstrate how entities that are present on the screen can be passed into an LLM using a novel textual representation that effectively summarizes the user’s screen while retaining relative spatial positions of these entities.
We show that ReaLM outperforms previous approaches, and performs roughly as well as the state-of-the-art LLM today, GPT-4, despite consisting of far fewer parameters, even for onscreen references despite being purely in the textual domain. It also outperforms GPT-4 for domain-specific user utterances, thus making ReaLM an ideal choice for a practical reference resolution system that can exist on-device without compromising on performance.
While our approach is effective in encoding the position of entities on the screen, we find that it results in loss of information that may not be able to resolve complex user queries that rely on nuanced positional understanding. We thus believe that exploring more complex approaches such as splitting the screen into a grid and encoding these relative spatial positions into text, while challenging, is a promising avenue of future exploration.
@@
While LLMs can generate unexpected output, our system offers the ability to constrain decoding or use simple post-processing to ensure this does not happen. Note however, that practically we find very little hallucination, and do not constrain the decoding of the LLM.
@@ Acknowledgements
The authors would like to thank Stephen Pulman, Leon Liyang Zhang, Jiarui Lu, Jeff Nichols, Shruti Bhargava, Dhivya Piraviperumal and Junhan Chen for their help and feedback.
@@ Appendix A Encoding onscreen entities
  • Onscreen Grab with Injected Turn Objects: Finally, the exact algorithm employed in our final approach is given in 2, and a sample encoding is shown in Table 7.
@@ 2       $surrounding\_objects\leftarrow$ Set of unique surrounding objects for $t$;
4       $clusters\leftarrow DBScan(surrounding\_objects,rect\_distance)$;
5
@@ 11                         Add $s$ to the prompt under key ‘surrounding_object‘;
Data: List of turn objects
Result: Onscreen parse constructed with turn object injection
1 $onscreen\_parse\leftarrow$ Empty list of onscreen parse elements;
2
// Step 0: Get all text boxes present in the screen
3 for each turn object $t$, index $i$ do
       // Step 1: Get unique surrounding objects
4       $surrounding\_objects\leftarrow$ Set of surrounding objects for $t$;
5
      // Step 2: Insert turn objects into the set (as turn objects are also text)
7
8
// Step 3: Sorting the centers of all surrounding objects
10
// Step 4: Determine vertical levels
11 $margin\leftarrow$ Margin for considering objects at the same level;
12 $levels\leftarrow$ List of vertical levels;
13 for each object $o$ in $sorted\_objects$ do
14       $same\_level\leftarrow$ List of objects at the same level as $o$;
15       for each object $other$ in $sorted\_objects$ do
17                   $same\_level\leftarrow same\_level\cup\{other\}$;
18
19
20      $levels\leftarrow levels\cup\{same\_level\}$;
21
22
// Step 5: Construct onscreen parse
23 for each level $l$ in $levels$ do
24       $level\_parse\leftarrow$ Empty string;
25       for each object $obj$ in $l$ do
[…]
27
[…]
29
30return $onscreen\_parse$;
Algorithm 2 Onscreen Parse Construction with Turn Object Injection
@@ Appendix C Sample Inputs
In this section, we show examples of how inputs into the model have been encoded, in the form of a visual representation.
@@
(a) Semantic Understanding
User Request: Call the evening Number
Screen:
{{1. 9 AM - 5 PM}}
{{2. 901.969.3120}}
{{3. 5 PM - 9 PM}}
{{4. 901.969.3391}}
Model Output: 4
(b) Summarisation
User Request: Remind me to get printouts before the tax deadline
Screen:
Tax Deadlines 2023
{{1. Feb 15}}
Reclaim your tax exemption from withholding
{{2. April 18}}
First-quarter estimated tax payment due
Model Output: 2
(c) World Understanding
User Request: Take me to the one in Washington
Screen:
Indian Embassy
{{1. 1701 El Camino Real, Mountain View 94040}}
{{2. 333 Dexter Ave N, Seattle 98109}}
{{3. 8295 Tournament Drive, Memphis, TN 38125}}
Model Output: 2
(d) Commonsense Reasoning
User Request: Save the link to the breakfast Recipe
Screen:
IMAGE
Strawberry Granola
{{1. Recipe link}}
IMAGE
Lavender boba tea
{{2. Recipe link}}
Model Output: 1
Figure 4: Qualitative Examples of LLM-based finetuned model able to adapt to complex use-cases

### v2

@@ 6.1 Analysis
GPT-4 $\approx$ ReALM $\gg$ MARRS for new use-cases: As a case study, we explore zero-shot performance of this model on an unseen domain: Alarms (we show a sample data point in Appendix Table 12). The last column in Table 3 compares the performance of all approaches and baselines on this unseen test set. We find that all of the LLM-based approaches outperform the FT model for this test set. Among the two, we find that the performance of ReALM and GPT-4 are very similar for the unseen domain. Additionally, Table 4 shows completely new experiences enabled by ReALM due to the LLM’s superior ability to perform complex understanding of natural language.
ReALM $>$ GPT-4 for domain-specific queries We find that due to finetuning on user requests, ReALM is able to understand more domain-specific questions. Consider Table 5. GPT-4 incorrectly assumes the reference to be about only a setting, whereas the ground truth consists of a home automation device in the background as well, and GPT-4 lacks the domain knowledge to be able to recognise that the device would also be relevant to this reference. ReALM, in contrast, doesn’t suffer from this due to being trained on domain-specific data.
@@ 7 Conclusion and Future Work
In this work, we demonstrate how large language models, which are typically trained on text alone, can be also be adapted to perform reference resolution to items in an extra-linguistic context. We do this by encoding entity candidates as natural text; we demonstrate how entities that are present on the screen can be passed into an LLM using a novel textual representation that effectively summarizes the user’s screen while retaining relative spatial positions of these entities. Our proposed system is thus able to resolve references in a variety of human-computer interaction settings, such as those involving on-screen, conversational and background entities; we note, however, that our proposed approach focuses primarily on anaphoric and deictic references, and we leave the extension of our system to handle other types of references, such as bridging references, to future work.
In addition, we show that ReALM outperforms previous approaches, and performs roughly as well as the current state-of-the-art LLM, GPT-4, despite consisting of far fewer parameters, even for onscreen references despite being purely in the textual domain. It also outperforms GPT-4 for domain-specific user queries, thus making ReALM an ideal choice for a practical reference resolution system that can exist on-device without compromising on performance.
While our approach is effective in encoding the position of entities on the screen, we find that it may not be able to resolve complex user queries that rely on nuanced positional understanding. We thus believe that exploring more complex approaches such as splitting the screen into a grid and encoding these relative spatial positions into text, while challenging, is a promising avenue of future exploration. In addition, in contrast to a critical assumption of our proposed system, not all on-screen entities are textual. While extending this paper to cover on-screen images, graphics and UI elements is beyond the scope of this work, this is certainly another extension that merits further investigation.
@@
While LLMs can generate unexpected output including potentially harmful text, our system offers the ability to constrain decoding or use simple post-processing to ensure this does not happen. Note however that practically we find very little hallucination or even text that deviates from the format that the models were finetuned on, and thus do not constrain the decoding of the LLM.
@@ Acknowledgements
The authors would like to thank Nidhi Rajshree, Stephen Pulman, Leon Liyang Zhang, Jiarui Lu, Jeff Nichols, Shruti Bhargava, Dhivya Piraviperumal, Junhan Chen and the anonymous reviewers for their help, suggestions, and feedback.
@@ Appendix A Encoding onscreen entities
  • Onscreen Grab with Injected Turn Objects: Finally, the exact algorithm employed in our final approach is given in 1, and a sample encoding is shown in Table 8.
@@ 2       $surrounding\_objects\leftarrow$ Set of unique surrounding objects for $t$;
4       $clusters\leftarrow DBScan(surrounding\_objects,$
5       $rect\_distance)$;
6
@@ Appendix C Sample Inputs
In this section, we show examples of how inputs into the model have been encoded, in the form of a visual representation in Tables 10, 11 and 12.